OpenStudy (anonymous):

The dimensions of a conical funnel are shown below. Nisha closes the nozzle of the funnel and fills it completely with a liquid. She then opens the nozzle. If the liquid drips at the rate of 15 cubic inches per minute, how long will it take for all the liquid to pass through the nozzle? (Use π=3.14.)

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

First you need to find the volume of the cone

jimthompson5910 (jim_thompson5910):

The volume of any cone with radius 'r' and height h is V = (pi*r^2*h)/3

OpenStudy (anonymous):

75.36

jimthompson5910 (jim_thompson5910):

Once you know the volume of the cone, you can use it to figure out how long that amount of liquid will take to drip out of the cone

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

Now we have this idea Amount dripped out = (Rate of drip)*(Time)

jimthompson5910 (jim_thompson5910):

So this means Amount dripped out = (Rate of drip)*(Time) 75.36 = 15*t where t is the time Note: we set "Amount dripped out" to all 75.36 cubic inches because want all the liquid to drip out

jimthompson5910 (jim_thompson5910):

So your last step is to solve for t

OpenStudy (anonymous):

5.024

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

so roughly 5 minutes (if you're rounding to the nearest minute)

jimthompson5910 (jim_thompson5910):

well nvm, you want to round up (to get that fraction portion too), so 6 minutes (if you're rounding)